# A Coreset Selection of Coreset Selection Literature: Introduction and Recent Advances

### v1

@@ I Introduction
Despite these early foundational works and comprehensive surveys and benchmarks, the literature now presents a wide range of coreset definitions and data reduction techniques that often differ significantly in formulation and scope [1, 15, 16, 17]. This divergence in definitions and methodologies highlights the evolution of the field and motivates the need for a unifying survey.
In contrast, this survey provides a structured and comprehensive synthesis of coreset selection strategies, categorizing them into three major paradigms: training-free, training-oriented, and blind coreset selection (i.e., approaches that select subsets without leveraging ground-truth labels). We discuss classical selection methods, such as random sampling and geometric approaches; training-informed techniques that leverage loss values, gradient norms, or decision boundaries; and emerging label-free strategies that use clustering or pre-trained models to estimate sample importance. Beyond algorithmic insights, we analyze the theoretical foundations of coreset selection and investigate how pruning strategies affect generalization, neural scaling laws, and dataset redundancy. Empirical comparisons highlight the trade-offs between computational efficiency and model performance across different methods.
Section II lays out definitions of coreset selection and active learning, a closely related field. The connection between these two areas is crucial because many coreset selection strategies naturally extend to active learning by selecting which samples to label. Section III presents training-free selection techniques, such as random sampling and geometry-based methods that still provide a strong baseline for more sophisticated methods. Section IV covers all approaches that fall under the umbrella term of training-oriented methods. They exploit information about training dynamics on the full dataset, such as scoring metrics, decision boundary estimations, submodular functions, gradient matching, and bilevel optimization. Next, Section V complements the previous sections by examining blind coreset selection (label-free data) with pseudo-labeling approaches or vision-language models. To provide practitioners with additional guidelines, Section VI provides more theoretical insights about coreset selection and when to apply which method. Section VII gives an overview of coreset applications to various fields in machine learning. Finally, this work summarizes and points to future directions in Section VIII and Section IX, respectively.
Figure 1: Venn diagram of coreset selection and active learning. Coreset selection typically assumes a labeled dataset, whereas active learning uses small labeled datasets and a large unlabeled pool. Many techniques are shared: both coreset selection and active learning rely on criteria (e.g., uncertainty, diversity, and coverage) to select data subsets.
@@ II Setting and Terminology
In this section, we formalize the coreset selection problem and highlight its close relationship to active learning, as shown in Figure 1. We also briefly touch upon theoretical perspectives that will guide deeper insights in later sections.
@@ II-A Problem Definition: Coreset Selection
We begin with the classic image classification setup, where the training dataset $\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ consists of $N$ i.i.d. samples drawn from an underlying data distribution $P$. We define $\mathbf{x}_{i}\in\mathcal{X}$ as the input, and $y_{i}\in\mathcal{Y}$ as its corresponding ground-truth label. Coreset selection aims to derive a subset $\mathcal{S}\subset\mathcal{T}$ with $|\mathcal{S}|\ll|\mathcal{T}|$ such that a model $\theta^{\mathcal{S}}$ trained on $\mathcal{S}$ generalizes as well as a model $\theta^{\mathcal{T}}$ trained on the full dataset $\mathcal{T}$:
@@
  • Training-Free: Methods that use geometric or statistical properties (e.g., distance in feature space or diversity measures) to select $\mathcal{S}$.
@@ II-B Problem Definition: Active Learning
A related research field is active learning, which also aims to reduce the required number of samples from a partially labeled perspective [22, 17, 23]. Instead of selecting a coreset $\mathcal{S}$ from a fully labeled dataset, active learning attempts to determine which unlabeled samples in a large pool $\mathcal{T}_{\mathcal{U}}$ should be sent to an oracle for labeling (e.g., a human annotator or nearest neighbor annotator) [24, 25, 26]. Thus, we aim to reduce the unlabeled dataset $\mathcal{T}_{\mathcal{U}}$, as illustrated in Figure 3.
Formally, consider a large unlabeled dataset $\mathcal{T}_{\mathcal{U}}$ and a small labeled dataset $\mathcal{T}_{\mathcal{L}}$, where $\mathcal{T}_{\mathcal{U}}\cap\mathcal{T}_{\mathcal{L}}=\emptyset$. The goal of active learning is to iteratively augment $\mathcal{T}_{\mathcal{L}}$ by selecting a subset $\mathcal{S}\subset\mathcal{T}_{\mathcal{U}}$ such that $|\mathcal{S}|\ll|\mathcal{T}_{\mathcal{U}}|$ while maximizing the expected model performance on $\mathcal{S}\cup\mathcal{T}_{\mathcal{L}}$. Active learning aims to solve the following optimization problem:
@@
where $\theta^{\mathcal{S}}$ is the model trained on $\mathcal{T}_{\mathcal{L}}\cup\mathcal{S}$, $\alpha\in(0,1)$ is the fraction of data left unlabeled, and $\mathcal{L}(\mathbf{x},y;\theta)$ is the model loss (e.g., cross-entropy loss).
In practice, several coreset strategies, especially those focusing on “most informative” or “diverse” examples, can be adapted to active learning settings by simply applying them to unlabeled data and then requesting annotations [27, 28]. Conversely, many established active learning techniques also serve as strong coreset baselines (if the samples are already labeled) [29, 22].
Figure 3: The active learning pipeline. The idea of active learning is to prune a large pool of unlabeled dataset $\mathcal{U}$ and to extend the already existing labeled dataset with essential, newly annotated data.
@@ II-C Coverage
[…]
@@ III-A Random Selection
Despite its simplicity, random sampling remains an important reference in coreset research for two reasons. First, its computational minimalism requires zero overhead beyond the selection step, i.e., no model training, no extra distance computations, etc. Second, in over-parameterized settings or for large, diverse datasets, even a random coreset can perform competitively [31, 15]. Redundant examples in $\mathcal{T}$ indicate that a randomly chosen $\mathcal{S}$ often covers a substantial fraction of the data distribution [32, 33].
@@ III-B Geometry Based
A fundamental and well-known assumption in machine learning is that samples closer in feature space tend to share similar properties [34, 35]. Geometry-based methods exploit this intuition by identifying redundant or noninformative samples while retaining a subset that effectively represents the original data distribution [36].
@@ III-B1 Herding
Herding is a deterministic coreset selection method that sequentially selects samples that best preserve the statistical properties of the full dataset [37, 38]. Unlike random sampling, Herding greedily selects points to minimize the discrepancy between the coreset mean and the dataset mean in feature space, called moment constraint [39, 40].
Specifically, given a feature mapping $\phi:\mathcal{X}\to\mathbb{R}^{d}$, the moment constraint requires that the empirical mean of the selected coreset approximates that of the full dataset:
@@
One benefit of Herding is that it is easy to implement if a suitable embedding $\phi(\cdot)$ is available. While it preserves the global statistics (the mean), its focus may overlook more subtle distribution patterns (like multi-modal clusters). Moreover, it requires iterative updates, which can be costly for very large $\mathcal{T}$.
@@
where $d(\mathbf{x}_{i},\mathbf{x}_{j})$ is a predefined distance metric (e.g., Euclidean). The goal is to select $k=(1-\alpha)\cdot{|\mathcal{T}|}$ points that serve as centers, ensuring that every remaining point is close to at least one of these centers. Since solving this problem exactly is NP-hard, an efficient greedy approximation is typically used [22].
@@ IV Training-Oriented Methods
As shown in Figure 5, geometry-based methods like Herding and k-Center Greedy heavily rely on the embedding quality of the underlying pre-trained feature extractor. One way to circumvent this problem is to train the model on the target dataset that needs to be pruned.
However, when training on the target dataset is required, one can use more sophisticated methods that also take into account the training dynamics, which fall under the umbrella term of training-oriented methods. By tailoring a subset $\mathcal{S}$ to the actual training dynamics, these methods can often achieve higher accuracy than their training-free counterparts, albeit at the expense of additional computation.
In this section, we present five major families of training-oriented coreset strategies: (i) scoring-based, (ii) decision boundary-based, (iii) submodularity-based, (iv) gradient matching-based, and (v) bilevel optimization-based methods. Throughout, we highlight how each method leverages partial or full training signals, and discuss typical trade-offs regarding efficiency versus precision.
Figure 6: Illustration of scoring-based methods. Usually, they employ a metric derived from multiple training iterations that quantifies the impact of a sample on the training quality (e.g., with forgetting counts, gradient norms, or sensitivity scores). The coreset is then selected by including the highest scores first. Note that the typical power law distribution of scores is only found for sufficiently trained and parameterized neural networks that are used for deriving the scores. Under-parameterized neural networks would lead to a more uniform distribution. Further details can be found in our theoretical remarks in Section VI).
@@ IV-A1 Forgetting
This metric can categorize samples as forgettable (i.e., experiencing one or more forgetting events) or unforgettable (i.e., never forgotten once learned). As a result, we can rank samples based on the number of forgetting events. Empirical studies suggest that unforgettable examples are often well-separated from the decision boundary, which typically indicates redundant information [48, 47, 49]. Forgettable examples, on the other hand, contribute more to refining the decision boundary and generalization. Thus, coreset selection methods based on forgetting take samples with the highest forgetting counts first.

### v2

@@ I Introduction
Despite early foundational work and several surveys and benchmarks, the field now spans a broad range of coreset definitions and reduction objectives that differ substantially in formulation, assumptions, and the signals they exploit [2, 43, 37, 123]. This divergence in definitions and methodologies highlights the evolution of the field and motivates the need for a unifying survey. As a result, existing reviews often cover only a slice of the landscape (e.g., geometric coresets or active learning), leaving it unclear how methods relate, where they overlap, and which assumptions drive their behavior. In particular, the same method can be described under different problem statements (coverage, informativeness, optimization).
To address this fragmentation, we introduce a taxonomy that unifies training-free, training-oriented, and label-free coreset selection under one view. Using this structure, we survey classical baselines (random, geometry-based), training-driven selectors (loss/gradient/boundary, submodular, gradient matching, bilevel), and emerging label-free approaches (clustering, foundation-model embeddings). Beyond algorithmic insights, we analyze the theoretical foundations of coreset selection and investigate how pruning strategies affect generalization, neural scaling laws, and dataset redundancy. Empirical comparisons highlight the trade-offs between computational efficiency and model performance across different methods.
Section II lays out definitions of coreset selection and active learning, a closely related field. The connection between these two areas is crucial because many coreset selection strategies naturally extend to active learning by selecting which samples to label. Section III presents training-free selection techniques, such as random sampling and geometry-based methods that still provide a strong baseline for more sophisticated methods. Section IV covers all approaches that fall under the umbrella term of training-oriented methods. They exploit information about training dynamics on the full dataset, such as scoring metrics, decision boundary estimations, submodular functions, gradient matching, and bilevel optimization. Next, Section V complements the previous sections by examining label-free coreset selection with pseudo-labeling approaches or vision-language models. To provide practitioners with additional guidelines, Section VI provides more theoretical insights about coreset selection and when to apply which method. Section VII gives an overview of coreset applications to various fields in machine learning. Finally, this work summarizes and points to future directions in Section VIII and Section IX, respectively.
Figure 1: Venn diagram of coreset selection and active learning. Coreset selection typically assumes a labeled dataset, whereas active learning uses small labeled datasets and a large label-free pool. Many techniques are shared: both coreset selection and active learning rely on criteria (e.g., uncertainty, diversity, and coverage) to select data subsets.
@@ II Setting and Terminology
In this section, we formalize the coreset selection problem and highlight its close relationship to active learning, as shown in Figure 1. In Table I, we list all typical notations used throughout the survey. We also briefly touch upon theoretical perspectives that will guide deeper insights in later sections. To avoid ambiguity across closely related data-reduction settings, Table II summarizes the key distinctions between coreset selection, active learning, and label-free variants in terms of supervision, outputs, and available training signals.
@@ II-A Problem Definition: Coreset Selection
We begin with the classic and standard image classification setup, where the training dataset $\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ consists of $N$ i.i.d. samples drawn from an underlying data distribution $P$. We define $\mathbf{x}_{i}\in\mathcal{X}$ as the input, and $y_{i}\in\mathcal{Y}$ as its corresponding ground-truth label. Coreset selection aims to derive a subset $\mathcal{S}\subset\mathcal{T}$ with $|\mathcal{S}|\ll|\mathcal{T}|$ such that a model $\theta^{\mathcal{S}}$ trained on $\mathcal{S}$ generalizes as well as a model $\theta^{\mathcal{T}}$ trained on the full dataset $\mathcal{T}$:
@@
  • Training-Free: Methods that use geometric or statistical properties (e.g., distance in feature space or diversity measures) from the samples alone to select $\mathcal{S}$.
@@
Most coreset works implicitly assume access to labels, either because they enforce class-wise budgets (for class balance) or because the selection directly exploits supervised signals such as per-sample loss, margins, or gradients. When labels are not used during selection, the setting is typically referred to as label-free (also called unsupervised/ blind): the subset is chosen purely from the inputs. This distinction is particularly important in scenarios where labels are expensive or unavailable.
@@
TABLE I: Notation used throughout the survey.
Symbol Meaning
$\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ Full labeled training dataset with $N$ samples.
$\mathcal{S}\subset\mathcal{T}$ Selected coreset (subset of training data).
$|\mathcal{S}|/|\mathcal{T}|\leq 1-\alpha$ Budget constraint; $\alpha\in(0,1)$ is the pruning ratio.
$\theta^{\mathcal{S}},\;\theta^{\mathcal{T}}$ Model parameters trained on $\mathcal{S}$ vs. on full $\mathcal{T}$.
$\mathcal{T}_{\mathcal{U}},\;\mathcal{T}_{\mathcal{L}}$ Label-free pool and labeled set in active learning / semi-supervised settings.
$\mathcal{Q}\subset\mathcal{T}_{\mathcal{U}}$ Query batch selected for annotation in active learning.
$B_{d}(\mathbf{x},r)$ Ball of radius $r$ centered at $\mathbf{x}$ under metric $d$.
$f:2^{V}\rightarrow\mathbb{R}$ Submodular set function defined on ground set $V$; optimized via greedy maximization.
$\chi_{t}(\mathbf{x},y),\;\chi_{t}^{*}(\mathbf{x},y)$ GraNd score (gradient norm) and EL2N approximation at epoch $t$.
$s(\mathbf{x},y)$ Sample sensitivity; worst-case relative loss contribution.
$\gamma_{j}$ Sample weight / step-size coefficient in gradient-matching methods (CRAIG/GradMatch).
$\lambda$ Regularization / trade-off scalar (e.g., in GradMatch/RETRIEVE objectives).
$\epsilon,\;\beta,\;\beta^{\prime}$ Tolerance threshold; scaling-law exponents for $\epsilon(\mathcal{T})\propto|\mathcal{T}|^{-\beta}$ and pruned $\epsilon(\mathcal{S})\propto|\mathcal{S}|^{-\beta^{\prime}}$.
@@ II-B Problem Definition: Active Learning
Building on the distinction between label-aware coreset selection (where labels are available and often directly exploited) and label-free selection (where selection is driven purely by inputs), a closely related research field is active learning, which targets data efficiency from a partially labeled perspective [122, 123, 86]. Instead of selecting a coreset $\mathcal{S}$ from a fully labeled dataset, active learning assumes access to a large unlabeled pool $\mathcal{T}_{\mathcal{U}}$ and asks which label-free samples should be queried from an oracle for annotation (e.g., a human annotator or nearest neighbor annotator) [33, 52, 167]. Thus, the goal is to reduce the label-free pool $\mathcal{T}_{\mathcal{U}}$ by selecting an informative query batch $\mathcal{Q}\subset\mathcal{T}_{\mathcal{U}}$, as illustrated in Figure 3.
Formally, consider a large label-free dataset $\mathcal{T}_{\mathcal{U}}$ and a small labeled dataset $\mathcal{T}_{\mathcal{L}}$, where $\mathcal{T}_{\mathcal{U}}\cap\mathcal{T}_{\mathcal{L}}=\emptyset$. The goal of active learning is to iteratively augment $\mathcal{T}_{\mathcal{L}}$ by selecting a subset $\mathcal{S}\subset\mathcal{T}_{\mathcal{U}}$ such that $|\mathcal{S}|\ll|\mathcal{T}_{\mathcal{U}}|$ while maximizing the expected model performance on $\mathcal{S}\cup\mathcal{T}_{\mathcal{L}}$. Active learning aims to solve the following optimization problem:
@@
where $\theta^{\mathcal{S}}$ is the model trained on $\mathcal{T}_{\mathcal{L}}\cup\mathcal{S}$, $\alpha\in(0,1)$ is the fraction of data left label-free, and $\mathcal{L}(\mathbf{x},y;\theta)$ is the model loss (e.g., cross-entropy loss).
In practice, several coreset strategies, especially those focusing on “most informative” or “diverse” examples, can be adapted to active learning settings by simply applying them to label-free data and then requesting annotations [34, 74]. Conversely, many established active learning techniques also serve as strong coreset baselines (if the samples are already labeled) [149, 122].
TABLE II: Boundaries between related data-reduction paradigms. We contrast supervision assumptions, outputs, training-signal access, and primary goals.
Paradigm Label assumption Output artifact Training access / signals Primary objective & typical methods
Coreset selection Fully labeled dataset $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{N}$ Subset of real examples $\mathcal{S}\subset\mathcal{T}$ (optionally weighted) May be training-free or uses (partial) training dynamics (loss, gradients, margins) Match performance of training on $\mathcal{T}$ at budget $(1-\alpha)$.
Methods: Random; geometry/coverage (Herding, $k$-Center); scoring (Forgetting, GraNd/EL2N); submodular (Facility Location/GraphCut); gradient matching (CRAIG/GradMatch); bilevel (GLISTER).
Active learning Large label-free pool; labels queried from an oracle Query batch $\mathcal{Q}\subset\mathcal{T}_{u}$ to label; iteratively expands $\mathcal{T}_{l}$ Iterative; relies on model uncertainty/disagreement + diversity criteria Maximize accuracy per annotation budget (label efficiency).
Methods: uncertainty (entropy/margin), diversity/coverage batches (coresets/submodular), disagreement/committee, boundary/adversarial-based querying.
label-free coreset selection No ground-truth labels available during selection Subset $\mathcal{S}\subset\{x_{i}\}_{i=1}^{N}$ (sometimes pseudo-labels/weights) No supervised training signal; uses representation geometry or self-supervised proxies Preserve downstream utility under compute/storage constraints without labels.
Methods: coverage/diversity in foundation-model feature space (CLIP/DINO + $k$-Center/submodular); pseudo-labeling + supervised selectors (ELFS-style); consistency/disagreement under augmentations; multimodal agreement signals.
Semi-supervised coreset selection Small labeled $\mathcal{T}_{l}$ + large label-free $\mathcal{T}_{u}$ Subset $\mathcal{S}\subset\mathcal{T}_{u}$ (or weighted label-free set) Uses labeled supervision sparsely + SSL losses (consistency/pseudo-labeling) Pick label-free points that best complement $\mathcal{T}_{l}$ for generalization.
Methods: bilevel/validation-driven (RETRIEVE), confidence filtering + diversity, hybrid submodular objectives.
Dataset distillation / condensation Typically labeled (or labels generated) Synthetic set $\tilde{\mathcal{S}}=\{(\tilde{x}_{j},\tilde{y}_{j})\}_{j=1}^{M}$, $M\ll N$ Heavy optimization; matches gradients/trajectories/distributions of training on $\mathcal{T}$ Replace real data with a compact synthetic training set.
Methods: gradient/trajectory matching (MTT), distribution matching, bilevel objectives, generative priors; coreset selection can be pre-/post-processing.
Figure 3: The active learning pipeline. The idea of active learning is to prune a large pool of label-free dataset $\mathcal{U}$ and to extend the already existing labeled dataset with essential, newly annotated data.
@@ II-C Coverage
$c=\int_{X}\mathbf{1}_{\mathrm{\cup_{\mathbf{x}\in\mathcal{S}}B_{d}(\mathbf{x},r)}}(\mathbf{x})d\mathbb{P}(\mathbf{x}),$ (3)
@@ III-A Random Selection
Despite its simplicity, random sampling remains an baseline reference in coreset research for two reasons. First, its computational minimalism requires zero overhead beyond the selection step, i.e., no model training, no extra distance computations, etc. Second, in over-parameterized settings or for large, diverse datasets, even a random coreset can perform competitively [44, 43]. Redundant examples in $\mathcal{T}$ indicate that a randomly chosen $\mathcal{S}$ often covers a substantial fraction of the data distribution [80, 14].
@@ III-B Geometry Based
A fundamental and well-known assumption in machine learning is that samples closer in feature space tend to share similar properties [18, 46]. Geometry-based methods exploit this intuition by identifying redundant or noninformative samples while selecting a subset that effectively represents the original data distribution [10].
@@ III-B1 Herding
Herding sequentially selects samples that best preserve the statistical properties of the full dataset [150, 24]. Unlike random sampling, Herding greedily selects points to minimize the discrepancy between the coreset mean and the dataset mean in feature space, called moment constraint [61, 85]. Specifically, given a feature mapping $\phi:\mathcal{X}\to\mathbb{R}^{d}$, the moment constraint requires that the empirical mean of the selected coreset approximates that of the full dataset:
@@
One benefit of Herding is that it is easy to implement if a suitable embedding $\phi(\cdot)$ is available. While it preserves the global statistics (the mean), its focus may overlook more subtle distribution patterns (e.g., multi-modal clusters). Moreover, it applies iterative updates, which can be costly for very large $\mathcal{T}$.
@@
where $d(\mathbf{x}_{i},\mathbf{x}_{j})$ is a predefined distance metric. The goal is to select $k=(1-\alpha)\cdot{|\mathcal{T}|}$ points that serve as centers, ensuring that every remaining point is close to at least one of these centers. Since solving this problem exactly is NP-hard, an efficient greedy approximation is typically used [122].
@@
III-C Embedding Space
As shown in Figure 5, geometry-based methods like Herding and k-Center Greedy heavily rely on the embedding quality of the underlying pre-trained feature extractor. A recent shift is to replace task-trained backbones (e.g., supervised ImageNet encoders) with foundation-model representations that are substantially more transferable and semantically aligned across domains [159, 41, 146]. In vision, strong options include large-scale self-supervised ViTs such as DINO/DINOv2 [19, 108] and vision-language pretraining such as CLIP-style models [115]. In practice, this turns many classical training-free selectors into plug-and-play methods: given an encoder $\phi$, we embed each image $\mathbf{z}_{i}=\phi(\mathbf{x}_{i})$ and run the usual selection objective (e.g., $k$-Center or Herding) directly in the representation space.
With stronger embeddings, distance/similarity becomes a better proxy for semantic redundancy and coverage. This perspective also unifies training-free selection with recent label-free works: they often differ mainly in which encoder is used (self-supervised vs. vision-language) and whether additional proxies (e.g., pseudo-labels or difficulty estimates) are layered on top.
@@ IV Training-Oriented Methods
One way to circumvent the embedding space problem is to train the model on the target dataset that needs to be pruned. However, when training on the target dataset is required, one can use more sophisticated methods that also take into account the training dynamics, which fall under the umbrella term of training-oriented methods.
By tailoring a subset $\mathcal{S}$ to the actual training dynamics, these methods can often achieve higher accuracy than their training-free counterparts, albeit at the expense of additional computation. In this section, we present five major families of training-oriented coreset strategies. These include (i) scoring-based, (ii) decision boundary-based, (iii) submodularity-based, (iv) gradient matching-based, and (v) bilevel optimization-based methods. Throughout, we highlight how each method leverages partial or full training signals.
Figure 6: Illustration of scoring-based methods. Usually, they employ a metric derived from multiple training iterations that quantifies the impact of a sample on the training quality (e.g., with forgetting counts, gradient norms, or sensitivity scores). The coreset is then selected by including the highest scores first. Note that the typical power law distribution of scores is only found for sufficiently trained and parameterized neural networks that are used for deriving the scores. Under-parameterized neural networks would lead to a more uniform distribution. Further details can be found in our theoretical remarks in Section VI.
In the following, we use importance as an umbrella term for a sample-specific utility proxy, a measurable signal that approximates the benefit of including $(\mathbf{x},y)$ under a fixed selection budget (typically in terms of generalization). Different paradigms instantiate this proxy differently: scoring-based methods estimate training influence via per-sample dynamics (loss/forgetting/gradient magnitude), submodular methods define set-level utilities that encourage coverage and non-redundancy, and gradient-matching/bilevel methods align selection with optimization or validation objectives. Thus, “importance” should be read as proxy-dependent, not as a single universal quantity.
@@ IV-A1 Forgetting
This metric can categorize samples as forgettable (i.e., experiencing one or more forgetting events) or unforgettable (i.e., never forgotten once learned). As a result, we can rank samples based on the number of forgetting events. Empirical studies suggest that unforgettable examples are often well-separated from the decision boundary [134, 139, 51]. This separation typically indicates redundancy. Forgettable examples, on the other hand, contribute more to refining the decision boundary and generalization. Thus, coreset selection methods based on forgetting take samples with the highest forgetting counts first.